1. A child starts sledding from rest, going down a 40 m long 7.5° incline, then coasting across a horizontal stretch. The mass of the sled + child is 35 kg, and the coefficient of kinetic friction is 0.060.
1
Expert's answer
2020-12-14T12:16:21-0500

Answer

Energy loss due to friction is given by

"E=\mu mgS cos\theta"

"E=0.060\times 35\times 9.8\times \cos 7.5°"

"E=20.40J"